OpenStudy (anonymous):

square root of 45 plus square root 45 plus square root 27

OpenStudy (anonymous):

(3squareroot5)sqaured X 3squareroot3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\sqrt{45}=\sqrt{9\times 5}=\sqrt{9}\times \sqrt{5}=3\sqrt{5}\]

OpenStudy (anonymous):

\[ \sqrt{45}+\sqrt{45}=2\sqrt{45}=2\times 3\sqrt{5}=6\sqrt{5}\]

OpenStudy (anonymous):

and \[\sqrt{27}=\sqrt{9\times 3}=3\sqrt{3}\] so your "final answer" is \[6\sqrt{5}+3\sqrt{3}\]